OpenStudy (anonymous):

0.7z+1.9+.1z=5.5-.4z

OpenStudy (mimi_x3):

0.8z +1.9 = 5.5-4z 4.8z = 3.6 z = 3.6/4.8

OpenStudy (mimi_x3):

aww saifoo's not here to help xD

OpenStudy (saifoo.khan):

IM HERE.

OpenStudy (mimi_x3):

LOL

OpenStudy (saifoo.khan):

the easy way, multiply the whole equal by 10.

OpenStudy (mimi_x3):

multiply ? wtf why ?

OpenStudy (anonymous):

so is it what mimi said whats the answer

OpenStudy (saifoo.khan):

0.7z+1.9+.1z=5.5-.4z \[10( 0.7z+1.9+.1z=5.5-.4z )\]\[7z+19+z = 55 -4z\]

OpenStudy (saifoo.khan):

i think it's easy to solve now. Isn't it?

OpenStudy (mimi_x3):

what's the point of extra steps ?

OpenStudy (saifoo.khan):

\[8z +19 =55-4z \]8z-4z = 55 - 19\[4z = 36 \]z = 9

OpenStudy (saifoo.khan):

That extra step helps u to solve the equation!

OpenStudy (saifoo.khan):

it makes that easy.

OpenStudy (mimi_x3):

yu srs ?! then im wrong argh

OpenStudy (saifoo.khan):

Yea u r! =(

OpenStudy (mimi_x3):

yeah *sigh* xD